Jaidev takes 2 whole 1/5 minutes to walk across the school ground. Rahul takes 7 / 4 minutes to do the 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Time taken by Jaidev to walk across the school ground =2 whole 1/5 = 11/5minutes
Time taken by Rahul to walk across the school ground = 7 / 4 minutes
Convert these fractions into like fractions
11 / 5 = 11 / 5× 4 / 4
= (11 × 4) / (5 × 4)
= 44 / 20
7 / 4 = 7 / 4 ×5 / 5
= (7 × 5) / (4 × 5)
= 35 / 20
Clearly, 44 / 20 > 35 / 20
11 / 5 > 7 / 4
∴Rahul takes less time than Jaidev to walk across the school ground
Difference = 11 / 5–7 / 4
= 44 / 20–35 / 20
= 9 / 20
Hence, Rahul walks across the school ground by 9 / 20 minutes

Thanks